OpenStudy (anonymous):

use implicit differentiation to find dy/dx if sinxy=5x+3y^2

OpenStudy (anonymous):

is that sin(xy)?

OpenStudy (anonymous):

Ooop, had to fix it: \[ \begin{array}{rcl} \sin (xy)&=&5x+3y^2 \\ \frac{d}{dx} [ \sin (xy) ] &=& \frac{d}{dx} [ 5x+3y^2 ] \\ \cos (xy) \left(y+x\frac{dy}{dx}\right) &=& 5+6y \frac{dy}{dx} \\ y \cos (xy) + x\frac{dy}{dx} \cos (xy)\ &=& 5+6y \frac{dy}{dx} \\ x\frac{dy}{dx} \cos (xy)\ &=& 5+6y \frac{dy}{dx} -y \cos (xy) \\ x\frac{dy}{dx} \cos (xy)\ -6y \frac{dy}{dx} &=& 5-y \cos (xy) \\ (x\cos (xy)\ -6y) \frac{dy}{dx} &=& 5-y \cos (xy) \\ \frac{dy}{dx} &=& \frac{5-y \cos (xy)}{x\cos (xy)\ -6y} \end{array} \]

OpenStudy (anonymous):

Thank you so much!